OpenStudy (anonymous):

Systems Using Substitution 3x + 2y = 20 y = x + 4

OpenStudy (anonymous):

look at your previous post of the same question

OpenStudy (anonymous):

solve the first equation for x then substitute that x in for the x in the second equation in the second equation you are solving for y

OpenStudy (anonymous):

after you find y go back to your first equation and plug in your y and solve for x

OpenStudy (anonymous):

y = x+4 so u but that in place of y so you have 3x 2(x+4)=20 then simply solve for x

OpenStudy (anonymous):

afs os simpler

OpenStudy (anonymous):

correction i left out a plus 3x+2(x+4)=20 solve for x then plug in to other one and u will get y